OpenStudy (anonymous):

(4/x) + (4/x^2 - 9) = (3/x - 3)

OpenStudy (anonymous):

@absurdism

OpenStudy (anonymous):

Okay...I've worked on this one for ages. I think the LCD is x(x + 3)(x - 3). Want me to type up my work?

OpenStudy (anonymous):

let's just go through it

OpenStudy (anonymous):

K. Following your lead.

OpenStudy (anonymous):

to make this easy i'm just gonna multiply everything by x^2 to get rid of those ugly variables in the denominator 4x + 4 - 9x^2 = 3x - 3x^2 -6x^2 + x + 4 = 0 can you factor that

OpenStudy (anonymous):

I have another almost just like this, so when you've walked me through this one maybe I'll be able to solve the second.

OpenStudy (anonymous):

Um..I'm not sure I know how to, because I don't know of any factors of -24 that sum to 1. Show me?

OpenStudy (anonymous):

Also, how did you get x^2?

OpenStudy (anonymous):

Or rather, know to use x^2.

OpenStudy (anonymous):

I don't mean to make you do my work for me, I just don't know where to start there. @absurdism

OpenStudy (anonymous):

well it was just logic i have a weird intuition for problems like this (for all of math really), so i do things in a weird way but it works if we multiply everything by x^2 then it gets rid of that one x^2 in the denominator and the other x's in the other denominators idk if you understand but it makes sense to me

OpenStudy (anonymous):

I get why it cancels the x^2 in the middle denominator, but I guess I thought it would create x^3 in the other two. I don't think I could repeat that method.

OpenStudy (anonymous):

nope, 4/x * x^2 = (4x^2)/x = 4 * x^2/x = 4x

OpenStudy (anonymous):

ew that looks so gross one sec

OpenStudy (anonymous):

$$\frac{4}{x}(x^2) = \frac{4x^2}{x} = 4 \frac{x^2}{x} = 4x$$

OpenStudy (anonymous):

*cries* @absurdism

OpenStudy (anonymous):

\[\dfrac{4}{x}+\dfrac{4}{x^2}-9=\dfrac{3}{x}-3\] Is this your problem??

OpenStudy (anonymous):

ya i could have the problem wrong

OpenStudy (anonymous):

I have to confirm first.

OpenStudy (anonymous):

No, Bebong. Absurdism was right. I had them in parentheses to avoid it looking like that.

OpenStudy (anonymous):

I get it now, absurdism. K...on to the next step?

OpenStudy (anonymous):

\[\dfrac{4}{x}+\dfrac{4}{x^2-9}=\dfrac{3}{x-3}\]

OpenStudy (anonymous):

Yes, that is correct @Bebong

OpenStudy (anonymous):

So for the next part, do we get 4/-9??

OpenStudy (phi):

I would clear the denominators. Multiply both sides by the LCD x(x-3)(x+3) \[ \dfrac{4}{x}+\dfrac{4}{x^2-9}=\dfrac{3}{x-3} \\ 4(x-3)(x+3) + 4x = 3x(x+3) \] I skipped some steps, but if you multiply each term by the LCD and simplify you get that result. Now expand \[ 4(x^2 -9) + 4x = 3x^2 +9x\\ 4x^2 -36 +4x = 3x^2 +9x \\ x^2 -5x -36=0 \]

OpenStudy (anonymous):

I get the right side of that...how did you keep 4(x^2 - 9) + 4x from becoming 4(x^2 - 9 + x)???

OpenStudy (phi):

do you see how I got the second line \[ 4(x-3)(x+3) + 4x = 3x(x+3) \]

OpenStudy (anonymous):

Whoohoo, that gives me a usable answer...it factors out to (x - 9)(x + 4), which gives me the solutions of x = 9, x = -4.

OpenStudy (anonymous):

Yes, I get the factoring - what I don't get is how you kept the individual x from the LCD out of the parentheses...

OpenStudy (anonymous):

Because when I multiply 4 by x(x - 3)(x + 3), I get 4(x^2 + x - 9). And that doesn't factor.

OpenStudy (phi):

ok, in slow motion \[ \cancel{x}(x - 3)(x + 3)\cdot \dfrac{4}{\cancel{x}}+x\cancel{(x - 3)(x + 3)}\cdot \dfrac{4}{\cancel{x^2-9}}=\dfrac{3}{\cancel{x-3}} \cdot x\cancel{(x - 3)}(x + 3) \] that gives \[ 4(x-3)(x+3) +4x = 3x(x+3) \] the first term is 4(x^2-9) = 4x^2-36

OpenStudy (phi):

**Because when I multiply 4 by x(x - 3)(x + 3), I get 4(x^2 + x - 9) remember the first term is 4/x so the x's cancel

OpenStudy (anonymous):

AHA!

OpenStudy (phi):

it's not a coincidence the denominators "go away". That is what happens when we multiply by the LCD

OpenStudy (anonymous):

I get it!! Thank you @absurdism @phi you guys rock :)

OpenStudy (anonymous):

Sorry I can't give you both medals. @phi can you help me with one more like this - mainly just watch to make sure I'm doing the steps right like you showed me?

OpenStudy (phi):

please make it a new post.

OpenStudy (anonymous):

Sure :) Thanks!

OpenStudy (anonymous):

sorry i couldn't explain as well as phi could. ^-^

OpenStudy (anonymous):

That's fine @absurdism you still helped a lot!